OpenStudy (anonymous):

write an equation in the slope intercept form for the line passing through points (-2,0) and (0,2)

OpenStudy (anonymous):

can you find the slope using these points???

OpenStudy (anonymous):

\[m=\frac{ y2-y1 }{ x2-x1 }\] y = mx +b

OpenStudy (anonymous):

can you take it from here

OpenStudy (anonymous):

Yes, thank you!! @surdawi

OpenStudy (anonymous):

you welcome; make sure you write your answer when you are done and i will check it for you

OpenStudy (anonymous):

y=-2x+1

OpenStudy (anonymous):

slope is wrong

OpenStudy (anonymous):

want to retry it?

OpenStudy (anonymous):

can you write the equation above with numbers?

OpenStudy (anonymous):

y=x+4?

OpenStudy (anonymous):

intercept is wrong

OpenStudy (anonymous):

slope right wrong b

OpenStudy (anonymous):

can you show me how you solve for b so i can tell you where you went wrong

OpenStudy (anonymous):

y=x+2

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

use point slope form to get value of b. y-y1=m(x-x1)

OpenStudy (anonymous):

finally! thanks

OpenStudy (anonymous):

did you know where you made your mistake

OpenStudy (anonymous):

not really... I am really lost.

OpenStudy (anonymous):

so did you just guessed about 4 or what

OpenStudy (anonymous):

No, I had both -2 and 2 on the x axis and counted my way frm -2 to 2 and got 4. But I was wrong bc it was supposed to be 2 on the y axis

OpenStudy (anonymous):

yes but using formula can eliminate ambiguities i told you before the point slope form y-y1=m(x-x1) using this formula you can easily get the slope intercept form of line.